Type 2: In a population of 2000 gaboon vipers, a genetic difference with respect to venom exists at a single locus. The alleles are incompletely dominant. The population shows 100 individuals homozygous for the ‘t’ allele (genotype “tt” = poisonous) 800 heterozygous (genotype Tt = mildly poisonous, and 1100 homozygous for eth T allele (genotype TT = lethally poisonous). (a) What is the frequency of the ‘t’ allele in the population? (b) Are the genotypes in Hardy-Weinberg equlibrium?
Answer #1

As per the data,

Total population of Gaboon vipers= 2000

population with "tt" or poisonous trait = 100

=> frequency of tt = 100/2000 = 0.05

frequency of "t" = \sqrt{t} = 0.2236

population with "TT" or poisonous trait = 1100

=> frequency of TT = 1100/2000 = 0.55

frequency of "t" = = 0.2236frequency of "T" =  = 0.7416

the frequency of the ‘T’ allele in the population= 07416

What is the frequency of the ‘t’ allele in the population= 0.55

equation is p+q = 1

0.74

as per the equation =( p+q) = 1 = 0.74+0.05